OpenStudy (anonymous):

Please help MEDAL and FAN!!!! To simplify 6y + 2y, a student showed the following steps: 6y + 2y = y * 2 Line 1 = y(6 + 2) Line 2 = y(8) = 8y Which properties of real numbers are used in these steps? A) the Commutative and Associative Properties B) the Commutative and Distributive Properties C) the Addition Property of Equality and the Associative Property D) the Associative and Distributive Properties

OpenStudy (anonymous):

http://library.thinkquest.org/2647/algebra/post.htm bottom

OpenStudy (anonymous):

I dont understand the way it explains that :/

OpenStudy (anonymous):

its b...Commutative and Distributive Properties

OpenStudy (anonymous):

how do you get it?

OpenStudy (anonymous):

6y + 2y = y * 2 Line 1 i don't get this line. is it saying 6y+2y=2y?

OpenStudy (anonymous):

I actually used what tomo said and I went to his website link and it had at the bottom the explanations for Commutative and Distributive Properties

OpenStudy (anonymous):

y times 2 @tomo

OpenStudy (anonymous):

i don't think that can be true. take for instance, 6*1 + 2*1 = 2*1 which is not true.

OpenStudy (anonymous):

its just what the question said....My math teacher makes all the questions, and hes not very bright when it comes to math...and he even said that

OpenStudy (anonymous):

GO with b :)

OpenStudy (anonymous):

ok thank you :)

OpenStudy (anonymous):

no problem :) have a good day (:

OpenStudy (anonymous):

you too

OpenStudy (anonymous):

:)